OpenStudy (anonymous):

If the level of a lake increases steadily by 8 inches over 2 years, what is the change in lake level during each year? negative sixteen inches negative four inches 4 inches sixteen inches

OpenStudy (anonymous):

Change = 8/2 = 4/1 or 4 inches per year

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

:)